# Perfect Secret Key Generation for a class of Hypergraphical Sources

### v1

@@ I Introduction
The multiparty secret key generation problem considers $m$ parties that need to agree upon a group secret key. To facilitate secret key generation, the parties are given access to a discrete memoryless multiterminal source (DMMS), i.e., i.i.d. copies of distinct but correlated random variables, and are allowed to communicate interactively over a noiseless but public broadcast channel. The secret key generated therefore, needs to be secured from any passive eavesdropper [^1] An eavesdropper is said to be passive if it can only listen to the messages but not tamper with them. listening to the communication sent over the channel. Initially posed for $m=2$ parties independently by [Maurer93, AC93], this problem was later extended to an arbitrary number of parties by the seminal work of Csiszár and Narayan [CN04].
The key quantity of interest in these works is the secret key capacity, i.e., the maximum rate of a secret key that can be generated, a full characterization of which appears in [CN04]. An alternate characterization of the secret key capacity was given by Chan and Zheng [CZ10] who showed that the secret key capacity equals a multivariate analogue of the mutual information [chan15], which we shall refer to as multipartite information.
@@
The PIN model is defined using an underlying graph whose vertex set equals the set of parties. Associated with every edge are i.i.d. sequences of independent Be(1/2)(1/2) random variables, and each party has access to the random variables associated with the edges incident with it. Nitanawarat and Narayan [nitinrawat] gave an elegant perfect secret key generation scheme for the PIN model achieving capacity. The scheme involves packing multiple copies of the underlying graph with spanning trees, and generating a single bit of perfect secret key from every spanning tree. This scheme therefore achieves a perfect secret key of rate equal to the spanning tree packing rate of the graph, which is shown to equal the (strong) secret key capacity by exploiting standard results on spanning tree packing such as Tutte’s Theorem [diestel, Theorem 2.4.1].
@@
Our contributions in this work are two-fold. Firstly, we provide an explicit secret key achieving scheme for the complete $t$-uniform hypergraph by exploiting its combinatorial properties. We begin by developing a key generation protocol that extracts $\binom{m-2}{t-2}$ bits of perfect secret key from star hypergraphs whose projected hypergraph is complete – see Section III for formal definitions. Therefore, any packing of the complete $t$-uniform hypergraph by such star hypergraphs leads to a perfect secret key of rate equal to the packing rate times $\binom{m-2}{t-2}$. The scheme is then shown to match the secret key capacity by constructing a packing of sufficient rate. We remark here that our scheme is closely related to the common randomness generation scheme for hypergraphical sources appearing in [HTKZPW].
@@ II-A Multiterminal Source Model
The multiterminal source model of Csiszár and Narayan [CN04] considers a set $\mathcal{M}=\{1,2,\dots,m\}$ of $m\geq 2$ parties. Party $i\in\mathcal{M}$ observes an $n$-length random vector $X_{i}^{n}$ taking values in a finite alphabet $\mathcal{X}_{i}^{n}$. The collection of observations across all parties is denoted by $X_{\mathcal{M}}^{n}=(X_{i}^{n}:i\in\mathcal{M})$. We assume the source is memoryless, meaning the joint distribution factors as $P_{X_{\mathcal{M}}^{n}}(x_{\mathcal{M}}^{n})=\prod_{t=1}^{n}P_{X_{\mathcal{M}}}(x_{\mathcal{M},t})$, where $P_{X_{\mathcal{M}}}$ is a joint probability mass function on $\prod_{i\in\mathcal{M}}\mathcal{X}_{i}$. For any subset $A\subseteq\mathcal{M}$, we define the notation $X_{A}=(X_{i}:i\in A)$.
@@
Of relevance to us is the partition of $\mathcal{M}$ into singletons, denoted by $\mathcal{S}=\{\{1\},\{2\},\ldots,\{m\}\}$, which we shall henceforth refer to as the singleton partition. Following the convention of [mukherjee], we shall refer to the sources where $\mathcal{S}$ is a minimizer of the right hand side of (1) as Type-$\mathcal{S}$ sources. Therefore, $C_{\mathrm{SK}}(X_{\mathcal{M}})=I_{\mathcal{S}}(X_{\mathcal{M}})$ for Type-$\mathcal{S}$ sources.
While $C_{\mathrm{SK}}(X_{\mathcal{M}})$ is defined under the asymptotic notion of strong secrecy (where error and leakage vanish as $n\to\infty$), there is a stricter non-asymptotic notion known as perfect secrecy. A random variable $K$ is a perfect secret key if it satisfies exact recoverability ($\Pr(K_{1}=\dots=K_{m}=K)=1$), perfect secrecy ($I(K;F)=0$), and perfect uniformity ($H(K)=\log|\mathcal{K}|$) for some fixed blocklength $n$. It is immediate that a perfect secret key satisfies strong secrecy, but not vice versa. Accordingly, the perfect secret key capacity is upper bounded by the usual strong secret key capacity.
We remark that henceforth, the phrase ‘secret key capacity’ will be used to refer to the strong secret key capacity. Furthermore, we shall say that a perfect secret key achieves capacity if it has a rate equal to the secret key capacity. This work shows that for instances of a special class of sources, called the hypergraphical sources, perfect secret keys achieve capacity.
@@ II-B Hypergraphical Source Model
A hypergraphical source is defined using an underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ whose vertex set is the set of parties $\mathcal{M}$. Associated with each hyperedge $e\in\mathcal{E}$ are $n$ i.i.d. copies of a Bernoulli(1/2) random variable $\xi_{e}$. Furthermore, the collection of random variables $(\xi_{e}:e\in\mathcal{E})$ is mutually independent. The observation of party $i\in\mathcal{M}$ is the collection of variables associated with hyperedges incident to it, i.e., $X_{i}=(\xi_{e}:i\in e)$. Consequently, the joint source is $X_{\mathcal{M}}=(\xi_{e}:e\in\mathcal{E})$.
@@
A special case of the hypergraphical source is the pairwise independent network (PIN) model [pinorig, nitinrawat], where the underlying hypergraph is a graph. Nitinawarat and Narayan [nitinrawat] has shown that the PIN model admits a linear communication scheme which achieves a linear perfect secret key, via spanning tree packing of the underlying graph. In this work, we develop similar linear perfect secret key generation schemes for $t$-uniform hypergraphical sources, by exploiting the combinatorial structure of the underlying hypergraph.
@@
We next show that any perfect omniscience scheme can be leveraged to generate a perfect secret key. To proceed, we need the following technical lemma.

### v2

@@ I Introduction
The multiparty secret key generation problem considers $m$ parties that need to agree upon a group secret key. To facilitate secret key generation, the parties are given access to a discrete memoryless multiterminal source (DMMS), i.e., i.i.d. copies of distinct but correlated random variables, and are allowed to communicate interactively over a noiseless but public broadcast channel. The secret key generated, therefore, needs to be secured from any passive eavesdropper [^1] An eavesdropper is said to be passive if it can only listen to the messages but not tamper with them. listening to the communication sent over the channel. Initially posed for $m=2$ parties independently by [Maurer93, AC93], this problem was later extended to an arbitrary number of parties by the seminal work of Csiszár and Narayan [CN04].
The key quantity of interest in these works is the secret key capacity, i.e., the maximum rate of a secret key that can be generated, a full characterization of which appears in [CN04]. An alternate characterization of the secret key capacity was given by Chan and Zheng [CZ10], who showed that the secret key capacity equals a multivariate analogue of the mutual information [chan15], which we shall refer to as multipartite information.
@@
The PIN model is defined using an underlying graph whose vertex set equals the set of parties. Associated with every edge are i.i.d. sequences of independent Be(1/2)(1/2) random variables, and each party has access to the random variables associated with the edges incident with it. Nitanawarat and Narayan [nitinrawat] gave an elegant capacity achieving perfect secret key generation scheme for the PIN model. The scheme involves packing multiple copies of the underlying graph with spanning trees and generating a single bit of perfect secret key from every spanning tree. This scheme therefore achieves a perfect secret key of rate equal to the spanning tree packing rate of the graph, which is shown to equal the (strong) secret key capacity by exploiting standard results on spanning tree packing such as Tutte’s Theorem [diestel, Theorem 2.4.1].
@@
Our contributions in this work are two-fold. Firstly, we provide an explicit secret key achieving scheme for the complete $t$-uniform hypergraph by exploiting its combinatorial properties. We begin by developing a key generation protocol that extracts $\binom{m-2}{t-2}$ bits of perfect secret key from star hypergraphs obtained from complete $t$-uniform hypergraphs – see Section III for formal definitions. Therefore, any packing of the complete $t$-uniform hypergraph by such star hypergraphs leads to a perfect secret key of rate equal to the packing rate times $\binom{m-2}{t-2}$. The scheme is then shown to match the secret key capacity by constructing a packing of sufficient rate. We remark here that our scheme is closely related to the common randomness generation scheme for hypergraphical sources appearing in [HTKZPW].
@@
The remainder of the paper is organised as follows. Section II-A introduces the problem setup, followed by Section II-B which details the key object of our study, the hypergraphical source model. The capacity achieving perfect secret key generation scheme for complete $t$-uniform hypergraphs appears in Section III. A perfect secret key generating scheme for $3$-uniform hypergraphs appears in Section IV. The paper concludes with Section V. Proofs of certain technical results are relegated to the Appendices for better readability.
@@ II-A Multiterminal source model
The multiterminal source model of Csiszár and Narayan [CN04] considers a set $\mathcal{M}=\{1,2,\dots,m\}$ of $m\geq 2$ parties. Party $i\in\mathcal{M}$ observes an $n$-length random vector $X_{i}^{n}$ taking values in a finite alphabet $\mathcal{X}_{i}^{n}$. The collection of observations across all parties is denoted by $X_{\mathcal{M}}^{n}=(X_{i}^{n}:i\in\mathcal{M})$. We assume that the source is memoryless, meaning that the joint distribution factors as $P_{X_{\mathcal{M}}^{n}}(x_{\mathcal{M}}^{n})=\prod_{t=1}^{n}P_{X_{\mathcal{M}}}(x_{\mathcal{M},t})$, where $P_{X_{\mathcal{M}}}$ is a joint probability mass function on $\prod_{i\in\mathcal{M}}\mathcal{X}_{i}$. For any subset $A\subseteq\mathcal{M}$, we define the notation $X_{A}=(X_{i}:i\in A)$.
@@
Of relevance to us is the partition of $\mathcal{M}$ into singletons, denoted by $\mathcal{S}=\{\{1\},\{2\},\ldots,\{m\}\}$, which we shall henceforth refer to as the singleton partition. Following the convention of [mukherjee], we shall refer to the sources where $\mathcal{S}$ is a minimizer of the right-hand side of (1) as Type-$\mathcal{S}$ sources. Therefore, $C_{\mathrm{SK}}(X_{\mathcal{M}})=I_{\mathcal{S}}(X_{\mathcal{M}})$ for Type-$\mathcal{S}$ sources.
While $C_{\mathrm{SK}}(X_{\mathcal{M}})$ is defined under the asymptotic notion of strong secrecy (where error and leakage vanish as $n\to\infty$), there is a stricter non-asymptotic notion known as perfect secrecy. A random variable $K$ is a perfect secret key if it satisfies exact recoverability ($\Pr(K_{1}=\dots=K_{m}=K)=1$), perfect secrecy ($I(K;F)=0$), and perfect uniformity ($H(K)=\log|\mathcal{K}|$) for some blocklength $n$. It is immediate that a perfect secret key satisfies strong secrecy, but not vice versa. Accordingly, the perfect secret key capacity is upper-bounded by the usual strong secret key capacity.
We remark that henceforth, the phrase ‘secret key capacity’ will be used to refer to the strong secret key capacity. Furthermore, we shall say that a perfect secret key achieves capacity if it has a rate equal to the secret key capacity. This work shows that, for instances of a special class of sources, called the hypergraphical sources, perfect secret keys achieve capacity.
@@ II-B Hypergraphical source model
A hypergraphical source is defined using an underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ whose vertex set is the set of parties $\mathcal{M}$. We allow $\mathcal{E}$ to be a multiset, i.e., $\mathcal{E}$ is allowed to have repeated elements. Associated with each hyperedge $e\in\mathcal{E}$ are $n$ i.i.d. copies of a Bernoulli(1/2) random variable $\xi_{e}$. Furthermore, the collection of random variables $(\xi_{e}:e\in\mathcal{E})$ is mutually independent. The observation of party $i\in\mathcal{M}$ is the collection of variables associated with hyperedges incident to it, i.e., $X_{i}=(\xi_{e}:i\in e)$. Consequently, the joint source is $X_{\mathcal{M}}=(\xi_{e}:e\in\mathcal{E})$.
@@
A special case of the hypergraphical source is the pairwise independent network (PIN) model [pinorig, nitinrawat], where the underlying hypergraph is a graph. Nitinawarat and Narayan [nitinrawat] have shown that the PIN model admits a linear communication scheme which achieves a linear perfect secret key, via spanning tree packing of the underlying graph. In this work, we develop similar linear perfect secret key generation schemes for $t$-uniform hypergraphical sources by exploiting the combinatorial structure of the underlying hypergraph.
@@
We next show that any perfect omniscience scheme can be leveraged to generate a perfect secret key.
{restatable}
lemmalinkey Let $X_{\mathcal{M}}$ be a hypergraphical source defined on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. For a blocklength $n\in\mathbb{N}$, let there exist a linear communication scheme of $r<n|\mathcal{E}|$ bits achieving perfect omniscience. Then there exists a perfect secret key of rate at least $|\mathcal{E}|-\frac{r}{n}$.
Proof:
See Appendix A. ∎
We conclude this section by showing how the notion of packing can be used in tandem with the perfect secret key generation scheme of Lemma II-B to increase the perfect secret key rate available from hypergraphical sources. A hypergraphical source $X_{\mathcal{M}}$ on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ is said to admit a packing for blocklength $n$, if there exists disjoint subsets $\mathcal{E}_{(1)},\mathcal{E}_{(2)},\ldots,\mathcal{E}_{(s)}\subseteq\mathcal{E}^{(n)}$. Then, $\mathcal{H}_{(l)}:=(\mathcal{M},\mathcal{E}_{(l)})$ can be viewed as a hypergraph in its own right. Moreover, we define $X_{i_{(l)}}:=(\xi_{e(j)}:i\in e,e(j)\in\mathcal{E}_{(l)})$, and consequently $X_{\mathcal{M}_{(l)}}:=(X_{i_{(l)}}:i\in\mathcal{M})$ is a hypergraphical source.
{restatable}
lemmapackedkeys Let $X_{\mathcal{M}}$ be a hypergraphical source on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. For a blocklength $n$, let $\mathcal{H}$ admit a packing, such that the hypergraphical source $X_{\mathcal{M}_{(l)}},1\leq l\leq s$, can generate a perfect secret key of $k_{l}$ bits using a blocklength of one. Then, $X_{\mathcal{M}}$ can generate a perfect secret key of rate $\frac{\sum_{l=1}^{s}k_{l}}{n}$.
Proof:
See Appendix A. ∎
III Perfect secret keys achieving capacity for complete $t$-uniform hypergraphs
This section considers hypergraphical sources on the complete $t$-uniform hypergraph $K_{m,t}$. The complete $t$-uniform hypergraph possesses every possible hyperedge of size $t$, i.e., $\mathcal{E}=\{e\subset\mathcal{M}:|e|=t\}$. $K_{m,t}$ is known to be Type-$\mathcal{S}$ [mukherjee], and this leads to a simple expression for its capacity.
@@ III Perfect secret keys achieving capacity for complete $t$-uniform hypergraphs
The secret key capacity of the hypergraphical source on $K_{m,t}$ is given by $\frac{t-1}{m-1}\binom{m}{t}$.
Proof:
This follows immediately from the fact that $K_{m,t}$ is Type-$\mathcal{S}$ [mukherjee, Corollary 19], the capacity expression for Type-$\mathcal{S}$ hypergraphical sources in (3), and noting $|\mathcal{E}|=\binom{m}{t}$.
To obtain the perfect secret key achieving capacity, we shall first obtain a packing of $K_{m,t}$ into what we refer to as star hypergraphs.
Definition 2 (Star Hypergraphs $S_{\mathcal{H},i}$).
Consider a hypergraphical source $X_{\mathcal{M}}$ defined on a hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$, and fix $i\in\mathcal{M}$. Then, the star hypergraph anchored at $i$ is the subhypergraph $S_{\mathcal{H},i}=(\mathcal{M},\mathcal{E}_{i})$. Furthermore, the hypergraphical source associated with $S_{\mathcal{H},i}$ is denoted by $X_{\mathcal{M}_{S_{i}}}:=(X_{1_{S_{i}}},\ldots,X_{m_{S_{i}}})$ where $X_{j_{S_{i}}}:=(\xi_{e}:i,j\in e)$, for all $j\in\mathcal{M}$.
For the remainder of this section, we shall use $S_{i}$ to denote $S_{K_{m,t},i}$ for notational convenience.
Our perfect secret key generation scheme for $K_{m,t}$ proceeds in two phases. We first show that for a blocklength $n=t$, the hypergraphical source on $K_{m,t}$ admits a packing into star hypergraphs. Then, we construct a perfect secret key generation scheme for each star hypergraph. Combining these results with Lemma II-B we are able to show that our key generation scheme gives perfect secret keys achieving capacity.
{restatable}
lemmasourcedecomp The hypergraphical source on $K_{m,t}$, for a blocklength $t$, admits a packing $\mathcal{E}_{(1)},\mathcal{E}_{(2)},\ldots,\mathcal{E}_{(m)}\subseteq\mathcal{E}^{t}$, where $K_{{m,t}_{(l)}}=(\mathcal{M},\mathcal{E}_{(l)}),1\leq l\leq m,$ is the star hypergraph $S_{l}$.
Proof:
See Appendix B. ∎
Next, we construct a linear communication achieving perfect omniscience for the source $X_{\mathcal{M}_{S_{i}}}$ on $S_{i}$.
Lemma 2.
For the source $X_{\mathcal{M}_{S_{i}}}$ on $S_{i}$, there exists a linear communication of $\binom{m-1}{t-1}-\binom{m-2}{t-2}$ bits achieving perfect omniscience. Hence, $S_{i}$ can generate a perfect secret key of rate $\binom{m-2}{t-2}$.
Proof:
We construct a communication scheme where only party $i$ sends messages. Fix any $j\neq i$. For any $e\in\mathcal{E}_{i}\setminus\mathcal{E}_{i,j}$, the party $i$ sends the message $(\oplus_{k\in e\setminus\{i\}}\xi_{(e\cup\{j\})\setminus\{k\}})\oplus\xi_{e}$. The communication scheme is linear by construction. Furthermore, the number of bits sent is $|\mathcal{E}_{i}\setminus\mathcal{E}_{i,j}|=\binom{m-1}{t-1}-\binom{m-2}{t-2}$. We first show that this communication scheme achieves perfect omniscience by showing that every $k\in\mathcal{M}$ recovers $\xi_{\mathcal{E}_{i}}$.
Obviously $i$ already knows $\xi_{\mathcal{E}_{i}}$. Next, note that for any $e\in\mathcal{E}_{i}\setminus\mathcal{E}_{i,j}$, the party $j$ can recover $\xi_{e}$ from the message $(\oplus_{k\in e\setminus\{i\}}\xi_{(e\cup\{j\})\setminus\{k\}})\oplus\xi_{e}$. Thus, $j$ perfectly recovers $\xi_{\mathcal{E}_{i}}$. Next, fix any $k\in\mathcal{M}\setminus\{i,j\}$. Since $j$ can recover $\xi_{\mathcal{E}_{i}}$ using the communication and $\xi_{\mathcal{E}_{i,j}}$, it is enough to show that $k$ can recover $\xi_{\mathcal{E}_{i,j}}$. Observe that $k$ already has access to $\xi_{e}$ for any $e\in\mathcal{E}_{i,j}$ such that $k\in e$. Now, fix any $e\in\mathcal{E}_{i,j}$ such that $k\notin e$. Consider the hyperedge $e^{\prime}=(e\cup\{k\})\setminus\{j\}\in\mathcal{E}_{i}\setminus\mathcal{E}_{i,j}$. Observe that $k$ has access to $\xi_{(e^{\prime}\cup\{j\})\setminus\{l\})}$ for every $l\in e^{\prime}\setminus\{i,k\}$. Then, $k$ can recover $\xi_{e}$ from the communication message $(\oplus_{l\in e^{\prime}\setminus\{i\}}\xi_{(e\cup\{j\})\setminus\{l\}})\oplus\xi_{e^{\prime}}$, as every hyperedge involved in this message contains $k$ except $e$.
Following perfect omniscience, the key generation scheme in Lemma II-B gives a perfect secret key of rate $|\mathcal{E}_{i}|-(\binom{m-1}{t-1}-\binom{m-2}{t-2})=\binom{m-2}{t-2}$. [^5] In fact, for $S_{i}$ we can explicitly define the perfect key to be $(\xi_{e}:e\in\mathcal{E}_{i,j})$. This key is obviously perfectly recoverable and uniform. The argument that it is perfectly secured from the communication follows by using a similar rank-based argument used in the proofs of Lemmas 10 and II-B.
We digress here to note that the perfect secret keys generated in Lemma 2 achieve capacity for $X_{\mathcal{M}_{S_{i}}}$. Using (3), we see that
$I_{\mathcal{S}}(X_{\mathcal{M}_{S_{i}}})=\frac{t-1}{m-1}|\mathcal{E}_{i}|=\frac{t-1}{m-1}\binom{m-1}{t-1}=\binom{m-2}{t-2}$.
An application of Lemma 1 immediately proves that $S_{i}$ is Type-$\mathcal{S}$ and the perfect secret key thus generated achieves capacity.
The packing scheme of Lemma 2 can now be combined with the key generation scheme for $S_{i}$ in Lemma 2 to generate perfect secret keys for $K_{m,t}$ achieving capacity.
Theorem 3.
There is an explicit perfect secret key generation scheme for $K_{m,t}$ that achieves capacity.
Proof:
Lemma 2 gives a packing $\mathcal{E}_{(1)},\ldots,\mathcal{E}_{(m)}\subseteq\mathcal{E}^{n}$ of $K_{m,t}$ for a blocklength $n=t$. Furthermore, $K_{{m,t}_{(i)}}=(\mathcal{M},\mathcal{E}_{(i)})$ is exactly the hypergraph $S_{i}$, for every $i\in\mathcal{M}$. Each of them can thus generate a perfect secret key of $\binom{m-2}{t-2}$ bits using the scheme of Lemma 2. Finally, by Lemma II-B, this generates a perfect secret key of rate
$\frac{m}{t}\binom{m-2}{t-2}=\frac{t-1}{m-1}\frac{m(m-1)}{t(t-1)}\binom{m-2}{t-2}=\frac{t-1}{m-1}\binom{m}{t}$, (4)
which by Lemma 1 equals the secret key capacity of $K_{m,t}$. ∎
IV Secret key generation scheme for $3$-uniform hypergraphs
In this section, we provide a new perfect secret key generation scheme for $3$-uniform hypergraphs (not necessarily complete), and we identify certain classes of 3-uniform hypergraphs where our scheme achieves capacity.
We begin by defining the notion of projection of a star hypergraph.
Definition 3.
Let $\mathcal{H}=(\mathcal{M},\mathcal{E})$ be a hypergraph, and let $i\in\mathcal{M}$. The projection of the star hypergraph $S_{\mathcal{H},i}$, denoted by $P_{\mathcal{H},i}$ is the hypergraph with vertex set $\mathcal{M}\setminus\{i\}$ and set of hyperedges $\{e\setminus\{i\}:e\in\mathcal{E}_{i}\}$.
The main ingredient in our construction is a type of $3$-uniform hypergraph which we call cycle-inducing hypergraph.
Definition 4.
A $3$-uniform hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ is said to be cycle-inducing if there exists $i\in\mathcal{M}$ such that $\mathcal{H}=S_{\mathcal{H},i}$ (i.e., $\mathcal{E}=\mathcal{E}_{i}$), and $P_{\mathcal{H},i}$ is a cycle [^6] For definitions of standard graph theory terms such as a cycle, refer to any standard textbook on graph theory such as [diestel]. on the vertex set $\mathcal{M}\setminus\{i\}$.
We call $i\in\mathcal{M}$ as the anchor.
The key result leading to our construction is the following proposition.
{restatable}
propositionprop Let $\mathcal{H}=(\mathcal{M},\mathcal{E})$ be a cycle-inducing $3$-uniform hypergraph with the anchor $i$. There exists a linear communication scheme of m3m-3 bits achieving perfect omniscience. As a consequence, one can generate a perfect secret key of 2 bits.
Proof:
See Appendix C. ∎
Before describing our generic scheme for $3$-uniform hypergraphs, we require one standard notion from graph theory, stated below.
Definition 5.
A (multi)graph GG [^7] Multigraph simply means that we allow the graph to have repeated edges. admits a Hamiltonian packing of size $r$ if there exists at least $r$ edge disjoint Hamiltonian cycles [^8] Recall that a cycle subgraph of a graph is called Hamiltonian if it includes all the vertices of the graph – see for example Chapter 10 of [diestel]. $C^{(1)},\ldots,C^{(r)}$ which are subgraphs of $G$.
If the set of cycles $C^{(1)},\ldots,C^{(r)}$ exhausts all the edges of the graph, then we say that $C^{(1)},\ldots,C^{(r)}$ is a Hamiltonian decomposition.
The following theorem now describes the perfect secret key generation scheme for $3$-uniform hypergraphs.
{restatable}
[…]
Proof:
See Appendix C. ∎
We now identify three different classes of $3$-uniform hypergraphs for which the scheme described in Theorem 9 is capacity achieving.
First, we consider the class of complete $3$-uniform hypergraph on $m$ vertices. {restatable}corollarytypea For the complete $3$-uniform hypergraph $K_{m,3}$ on $m$ vertices, the scheme of Theorem 9 with the set $A=\mathcal{M}$ achieves capacity.
Proof:
See Appendix C. ∎
Next, we consider a family of $3$-uniform hypergraphs $\mathcal{H}=(\mathcal{M},\mathcal{E})$ for which there exists $i\in\mathcal{M}$ such that $\mathcal{E}_{i}=\mathcal{E}$ and the graph $P_{\mathcal{H},i}$ admits a Hamiltonian decomposition. An interesting class of graphs admitting a Hamiltonian decomposition are Paley graphs [ALSPACH2012113]. More such classes can be found in [ALSPACH2012113].
{restatable}
corollarytypeb Consider a hypergraphical source on a $3$-uniform $\mathcal{H}=(\mathcal{M},\mathcal{E})$ for which there exists a vertex $i\in\mathcal{M}$ satisfying $\mathcal{E}_{i}=\mathcal{E}$ and the graph $P_{\mathcal{H},i}$ admits a Hamiltonian decomposition. Then, the perfect secret key generation scheme of Theorem 9 is capacity achieving. Furthermore, such a hypergraphical source is Type-$\mathcal{S}$.
Proof:
See Appendix C. ∎
We conclude our set of examples of $3$-uniform hypergraphs where the scheme of Theorem 9 achieves capacity with a class of hypergraphs we term Hollow 3D Kite hypergraphs. A Hollow 3D Kite hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ is defined as follows. The number of vertices $m=4r+1$ for some $r\geq 1$, and let $V_{1}=\{1,\dots,2r+1\}$ and $V_{2}=\{2r+2,\dots,4r+1\}$. The set of hyperedges is then defined as $\mathcal{E}=\left\{\{i,j,k\}:\{i,j\}\subseteq V_{1},k\in V_{2}\right\}$.
{restatable}
corollarytypec Consider a hypergraphical source on a $3$-uniform hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ which is a Hollow 3D Kite on $m=4r+1$ vertices. Then, the scheme of Theorem 9 gives perfect secret keys achieving capacity. Furthermore, the source is Type-$\mathcal{S}$.
Proof:
See Appendix C. ∎
V Conclusion
In this work, we gave explicit capacity achieving perfect secret key generation schemes for the complete $t$-uniform hypergraph and certain classes of $3$-uniform hypergraphs. Our schemes are an attempt to extend the spanning tree packing scheme of Nitinawarat and Narayan [nitinrawat] meant for PIN models on graphs. Unfortunately, there is no unique generalization of spanning trees to the setting of hypergraphs. Hence, for our scheme for $K_{m,t}$, we pack the hypergraph with star hypergraphs, which serve as the proxy for the spanning trees, and generate perfect secret keys for each star hypergraph. Similarly, the scheme for $3$-uniform hypergraphs uses cycle inducing $3$-uniform hypergraphs as stand-ins for spanning trees. We believe that a good generalization of spanning trees will shed further light on the problem of explicit perfect secret key generation schemes for hypergraphical sources.
In this regard, we believe that the so-called minimally topologically connected hypergraphs studied in [HTKZPW] are a good candidate for generalizations of spanning trees. Indeed, the common randomness generation scheme for minimally topologically connected hypergraphs studied in [HTKZPW] coupled with Lemma II-B gives an explicit construction of perfect secret keys achieving capacity. We conjecture that every topologically connected hypergraph can be suitably packed by minimally topologically connected hypergraphs such that invoking Lemma II-B will give a perfect secret key achieving capacity.
Furthermore, note that we have a scheme for $3$-uniform hypergraphs which achieves capacity for non-trivial classes of hypergraphs other than the complete hypergraph. A parallel direction of research would be to extend our scheme to $t$-uniform hypergraphs. This would require a generalization of the cycle to hypergraphs, and we conjecture that sphere triangulations is the correct generalization to look at in this case.
Appendix A Omitted proofs of Section II-B
To prove Lemma II-B, we first need the following technical lemma.
Lemma 4.